OpenStudy (anonymous):

Which solid has six faces, four lateral faces, two bases, eight vertices, and 12 edges?

Directrix (directrix):

Rectangular Solid

Directrix (directrix):

as known as rectangular prism

OpenStudy (anonymous):

For some reason I kept thinking it was a triangular prism x_x

Directrix (directrix):

I think the triangular prism has 9 edges. V + F = E + 2 for polyhedra where V = number of vertices, F = number of Faces, and E = number of edges. Note that a base is considered as a face.

OpenStudy (anonymous):

Ohh ! Thanks